OpenStudy (logan13):

Dennis is entering a contest to guess the number of marbles inside a cylinder. He wants to find the volume of the marbles and the volume of the cylinder to estimate how many marbles might be in the cylinder. Which formulas does he need to make a good guess? Choose all answers that are correct. (In these formulas, r represents radius, h represents height, l represents length, and w represents width.)

OpenStudy (logan13):

http://prntscr.com/ae1vcm

OpenStudy (logan13):

@Mehek14

OpenStudy (logan13):

@Zelda

OpenStudy (shadowlegendx):

He needs to find the volume of the cylinder, and the volume of the marbles So the first formula he needs is one to solve for the volume of a cylinder, whilst the second formula needs to be one that can solve for the volume for a sphere

OpenStudy (shadowlegendx):

Volume of a Cylinder: \[V = \pi r^2 h\] Volume of a Sphere \[V = \frac{ 4 }{ 3 } \pi r^2 h\]

OpenStudy (shadowlegendx):

The volume for a cylinder is there, so thatʻs one, but there isnʻt one that solves for that of a sphere, for the marbles

OpenStudy (shadowlegendx):

Dennis could however, use the formula \[V = \frac{ 1 }{ 3 } \pi r^2 h\] Then just multiply that by 4

OpenStudy (shadowlegendx):

That would suffice for the spheres

OpenStudy (logan13):

so A D